# A Scalable Optimization Approach for Equitable Facility Location: Methodology and Transportation Applications

### v3

@@ Abstract
Achieving equitable access to essential services, such as healthcare, food, and education, is a critical goal in urban planning and public policy. However, existing facility location models primarily focus on optimizing average accessibility, often neglecting equity concerns, particularly for disadvantaged populations. This paper proposes a novel, scalable framework for equitable facility location, introducing a linearized proxy for the Kolm-Pollak Equally-Distributed Equivalent (EDE) metric to balance efficiency and fairness. Extensive computational experiments demonstrate that, unlike other facility location models that incorporate measures of equity, our approach scales to extremely large problem instances. Optimal solutions represent significant improvements for the worst-off residents in terms of distance to an open amenity, while also attaining a near-optimal average experience for all users. Two real-world case studies—supermarket access and polling location access—illustrate the practical applicability of the framework. Furthermore, the model is extended to handle real-world considerations such as capacity constraints, split demand assignments, and location-specific penalties. By bridging the gap between equity theory and practical optimization, this work offers a robust and versatile tool for researchers and practitioners in urban planning, transportation, and public policy.
@@
location , equitable facility location , Kolm-Pollak Equally-Distributed Equivalent , integer programming , urban planning optimization
@@ 1 Introduction
The distribution of facilities within transportation networks shapes how resources and burdens are allocated across communities. From transit stops to emergency services, and from distribution centers to freight hubs, facility location decisions shape distributional justice. Traditional facility location models were designed to address operational problems and, as a consequence, typically provide solutions that are not concerned with the equity of the optimal distribution of distances (Current et al., 2002). Placing facilities to minimize mean distance tends to leave some community members many miles from service, while minimizing maximum distance ignores the average experience of the community.
@@
Despite these challenges, incorporating equity into facility location models is vital because it can help ensure that resources and services are distributed fairly across a population, reducing disparities and promoting social justice. This importance has led to a wealth of literature on the subject since the 1970s (O’Brien, 1969; McAllister, 1976; Savas, 1978). While various approaches have been proposed to incorporate equity in facility location models, including lexicographical optimization of multiple objectives (Ogryczak, 2009), these existing methods either do not scale computationally to city-sized instances, cannot evaluate both desirable and undesirable quantities, or lose their normative significance when applied to burdens rather than benefits (Barbati and Bruno, 2017). This has resulted in numerous proposals for incorporating equity considerations (Marsh and Schilling, 1994; Eiselt and Laporte, 1995; Drezner et al., 2009; Ogryczak, 2009; Lejeune and Prasad, 2013; Karsu and Morton, 2015; Barbati and Piccolo, 2016), but without consensus on an effective and tractable approach.
@@
The environmental justice literature has recently established clear criteria for metrics that can effectively evaluate distributions of both amenities and burdens (Logan et al., 2021). These properties include the ability to analyze population subgroups, consideration of both relative and absolute differences, and satisfaction of the mirror property - allowing consistent evaluation of both goods and their complementary burdens. The Kolm-Pollak Equally-Distributed Equivalent (EDE) has emerged as the preferred metric for environmental justice applications as it uniquely satisfies all these properties while providing a normative basis for comparing distributions (Maguire and Sheriff, 2020). However, its non-linear form makes it challenging to optimize, limiting its use in prescriptive facility location models that could help address inequities. This raises a critical question: How can we optimize the Kolm-Pollak EDE in facility location problems while maintaining computational tractability for real-world networks?
@@
We present a solution through a computationally tractable approach to optimize the Kolm-Pollak EDE in facility location problems. We prove equivalence between optimizing the non-linear EDE and a linear proxy, enabling solution of large real-world instances. We extend this formulation to handle split demands and capacity constraints common in transportation applications, and provide guidance for effectively scaling the inequality aversion parameter to achieve desired outcomes. We also develop and analyze an approach for incorporating location-specific penalties, providing theoretical bounds on approximation errors and practical guidelines for implementation. Through computational experiments on instances from the largest U.S. cities, including New York City with over 200 million binary variables, we demonstrate that our approach scales while producing solutions that balance average access with protection of worst-off residents.
@@
While we illustrate our methodology using supermarket access as an example, the approach is broadly applicable to transportation facility location problems including transit stop location, emergency service coverage, and distribution network design. Moreover, the mathematical framework developed for optimizing the Kolm-Pollak EDE could prove valuable for other domains where equity-based optimization of the distribution of resources or burdens is required.
@@ 2.1.1 Traditional Approaches and Their Limitations
    A set of residential areas $R$ (the ”origins” or ”customers”)
    A set of potential facility locations $S$ (points of interest)
    A target number of facilities $k$ to be opened
@@
    The population $p_{r}$ living there
    The distance $d_{r,s}$ to each potential facility location $s\in S$
@@
$x_{s}$ :=$:=\begin{cases}1&\text{if we open a facility at location }s\\ 0&\text{otherwise,}\end{cases}$ $\forall s\in S$
$y_{r,s}$ :=$:=\begin{cases}1&\text{if we assign residents from area }r\text{ to facility }s\\ 0&\text{otherwise,}\end{cases}$ $\forall s\in S,\forall r\in R$
@@
The p-median formulation minimizes the average distance traveled by residents to their assigned facility. The constraints ensure that: we open exactly $k$ facilities (1b); residents can only be assigned to open facilities (1c); and every residential area is assigned to exactly one facility (1d). However, by focusing solely on minimizing average distance, this approach can leave some community members unacceptably far from service, highlighting the need for models that explicitly consider equity.
2.1.2 Previous Attempts at Incorporating Equity
Early attempts to address equity in facility location focused on protecting the worst-off residents. The p-center model (Hakimi, 1965) takes this approach by minimizing the maximum distance that any resident must travel. This can be achieved by introducing a single variable $z$ that captures the maximum distance:
@@ 2.1.2 Previous Attempts at Incorporating Equity
Alternative approaches have been proposed to incorporate equity in facility location models. These include lexicographical optimization of multiple objectives (Ogryczak, 2009) and general p-norm optimization approaches (Karsu and Morton, 2015; Olivier et al., 2022; Gupta et al., 2023; Xinying Chen and Hooker, 2023). However, these existing methods either do not scale computationally to city-sized instances, cannot evaluate both desirable and undesirable quantities, or lose their normative significance when applied to burdens rather than benefits (Barbati and Bruno, 2017).
A recent computational study of eleven different equity measures found that several, including the log variance, Theil’s entropy coefficient, and the Atkinson equally-distributed equivalent, are too complex to incorporate into combinatorial optimization models (Barbati and Bruno, 2017). This has resulted in numerous proposals for incorporating equity considerations (Marsh and Schilling, 1994; Eiselt and Laporte, 1995; Drezner et al., 2009; Ogryczak, 2009; Lejeune and Prasad, 2013; Karsu and Morton, 2015; Barbati and Piccolo, 2016), but without consensus on an effective and tractable approach. To address these limitations, we require a measure that both appropriately captures equity concerns and remains computationally tractable for optimization.
@@ 2.2.1 Properties Required for Equity Assessment
The environmental justice literature has established clear criteria for metrics used to evaluate distributions of both amenities and burdens (Logan et al., 2021). There are six key properties that an inequality measure should possess if it is used to evaluate the distribution of a singular amenity or burden. For completeness, we note that a seventh and final property is included, however, this “multivariate” property is not relevant to our discussion and is omitted from the discussion below.
@@
2.2.2 Review of Existing Measures
The location of services and amenities has significant equity implications, as travel distance (for example, to grocery stores) represents a burden that can substantially impact community well-being, particularly for marginalized populations (Neff et al., 2009). Various metrics have been proposed to incorporate equity considerations in facility location, though each has important limitations.
@@ 2.2.2 Review of Existing Measures
The maximum value metric, commonly used in p-center models, focuses on optimizing the experience of the worst-off community member. However, both the maximum value and range metrics suffer from considering only extreme values while ignoring the overall distribution. Moreover, neither satisfies the principle of transfers - a key requirement for equity metrics. Several measures attempt to capture dispersion by examining deviation from the mean, including mean absolute deviation, variance, and variance of logarithms. However, these also fail to satisfy the transfers principle (Allison, 1978; Erkut, 1993).
@@
More sophisticated inequality measures have also been explored, including Schutz’s Index, the coefficient of variation, the Gini coefficient, Theil’s entropy coefficient, and Atkinson’s coefficient. While some of these metrics satisfy the transfers principle, they have other critical drawbacks. The Gini coefficient, despite its widespread use in facility location models (Mandell, 1991; Mulligan, 1991; Drezner et al., 2009; Lejeune and Prasad, 2013; Alem et al., 2022), measures only dispersion without considering the overall level of the distribution. This means it could prefer a solution where all residents are equally far from facilities over one with shorter but somewhat varied distances - clearly problematic for accessibility. The coefficient of variation and Theil’s coefficient share this limitation of ignoring scale.
@@
Equally-distributed equivalents (EDEs) offer a promising direction as they consider both the center and spread of distributions. The Atkinson EDE (Atkinson, 1970) has been applied to facility location, but it assumes higher values represent better welfare - making it inappropriate for measuring burdens like travel distance where smaller values are preferable (Cox Jr., 2012). This is because it does not satisfy the mirror property. Furthermore, computational studies have shown that optimizing the Atkinson EDE becomes intractable for realistically-sized problems (Barbati and Bruno, 2017).
@@
Indeed, computational tractability emerges as a major barrier for many proposed equity metrics. Recent comprehensive studies found that several measures including log variance, Theil’s coefficient, and the Atkinson EDE are too complex to incorporate into optimization models that can handle real-world problem sizes (Barbati and Bruno, 2017; Xinying Chen and Hooker, 2023). This computational challenge, combined with the theoretical limitations discussed above, has left the field without consensus on an appropriate and practical approach to measuring equity in facility location problems.
@@
These limitations motivated the development of the Kolm-Pollak EDE, which uniquely satisfies all key properties required for equity assessment in environmental justice applications. However, like many previous measures, its nonlinear form has limited its use to descriptive analysis rather than prescriptive optimization. This research bridges this gap - by developing a computationally tractable approach to optimizing the Kolm-Pollak EDE, we enable its application to large-scale facility location problems for the first time, allowing transportation planners to not just measure inequity but actively design for equity.

### v4

@@ Abstract
Efficient and equitable access to essential services, such as healthcare, food, and education, is an important goal in urban planning, public policy, and transport logistics. However, existing facility location models often do not scale well to large instances, or primarily focus on optimizing average accessibility, neglecting equity concerns, particularly for disadvantaged populations. This paper proposes a novel, scalable framework for equitable facility location, introducing a linearized proxy for the Kolm-Pollak Equally-Distributed Equivalent (EDE) metric to balance efficiency and fairness. Computational experiments demonstrate that our approach scales to extremely large problem instances, while being sensitive enough to account for inequity throughout the distribution, not merely via the maximum value. Moreover, optimal solutions represent significant improvements for the worst-off residents in terms of distance to an open amenity, while also attaining a near-optimal average experience for all users. An extensive real-world case study on supermarket access illustrates the practical applicability of the framework, with additional examples coming from polling applications. As such, the model is extended to handle real-world considerations such as capacity constraints, split demand assignments, and location-specific penalties. By bridging the gap between equity theory and practical optimization, this work offers a robust and versatile tool for researchers and practitioners in urban planning, transportation, and public policy.
@@
location , equitable facility location , Kolm-Pollak Equally-Distributed Equivalent , integer programming , urban planning optimization , discrete ordered median problem
@@ 1 Introduction
The distribution of facilities within transportation networks shapes how resources and burdens are allocated across communities. From transit stops to emergency services, and from distribution centers to freight hubs, facility location decisions shape distributional justice. Traditional facility location models were designed to address operational problems and, as a consequence, typically provide solutions that are not concerned with the equity of the optimal distribution of distances (Current et al., 2002). For example, placing facilities to minimize the mean distance tends to leave some community members many miles from service. On the other hand, prioritizing only the worst case, i.e., minimizing the maximum distance, ignores the average experience of the community.
@@
Despite these challenges, incorporating equity into facility location models is vital because it can help ensure that resources and services are distributed fairly across a population, reducing disparities, and promoting social justice. This importance has led to a wealth of literature on the subject since the 1970s (O’Brien, 1969; McAllister, 1976; Savas, 1978). Although various approaches have been proposed to incorporate equity in facility location models, including lexicographical optimization of multiple objectives (Ogryczak, 2009), these existing methods either do not scale computationally to city-sized instances, cannot evaluate both desirable and undesirable quantities, or lose their normative significance when applied to burdens rather than benefits (Barbati and Bruno, 2017). This has resulted in numerous proposals for incorporating equity considerations (Marsh and Schilling, 1994; Eiselt and Laporte, 1995; Drezner et al., 2009; Ogryczak, 2009; Lejeune and Prasad, 2013; Karsu and Morton, 2015; Barbati and Piccolo, 2016), but without consensus on an effective and tractable approach.
@@
The environmental justice literature has recently established clear criteria for metrics that can effectively evaluate distributions of amenities, such as clean air, and burdens, such as distance to an amenity, across a population (Logan et al., 2021). These (six) properties include the ability to analyze population subgroups, consideration of both relative and absolute differences, and satisfaction of the mirror property – allowing a consistent evaluation of both goods and their complementary burdens. The Kolm-Pollak Equally-Distributed Equivalent (EDE) has emerged as the preferred metric for environmental justice applications, as it uniquely satisfies all these properties while providing a normative basis for comparing distributions (Maguire and Sheriff, 2020). However, its non-linear form makes it challenging to optimize, limiting its use in prescriptive facility location models that could help address inequities. This raises a critical question: How can we optimize the Kolm-Pollak EDE in facility location problems while maintaining computational tractability for real-world networks?
@@
We present a solution by providing a computationally tractable approach to optimize the Kolm-Pollak EDE in facility location problems. We prove equivalence between optimizing the non-linear EDE and a linear proxy, enabling the solution of large real-world instances. We extend this formulation to handle split demands and capacity constraints common in transportation applications, and provide guidance for effectively scaling the inequality aversion parameter to achieve desired outcomes. We also develop and analyze an approach for incorporating location-specific penalties, providing theoretical bounds on approximation errors and practical guidelines for implementation. Through computational experiments on instances from the largest U.S. cities, including New York City with over 200 million binary variables, we demonstrate that our approach scales while producing solutions that balance average access with protection of worst-off residents.
@@
While we illustrate our methodology using supermarket access, this is only one example. Our approach is broadly applicable to transportation facility location problems, including transit stop location, emergency service coverage, and distribution network design. Moreover, the mathematical framework developed for optimizing the Kolm-Pollak EDE could prove valuable for any domain where equity-based optimization of the distribution of resources or burdens is required.
@@ 2.1 Facility Location Models
    a set of residential areas, $R$ (the “origins” or “customers”);
    a set of potential facility locations, $S$ (points of interest);
    a target number of facilities to be opened, $k$.
@@
    the population living there, $p_{r}$.
    the distance to each potential facility location, $d_{r,s}$, for all $s\in S$.
@@
$x_{s}$ $:=\begin{cases}1&\text{if we open a facility at location }s,\\ 0&\text{otherwise,}\end{cases}$ $\forall~s\in S;$
[…]
@@
The p-median formulation minimizes the average (or equivalently, total) distance traveled by residents to their assigned facility. The constraints ensure that: we open exactly $k$ facilities (1b); residents can only be assigned to open facilities (1c); and every residential area is assigned to exactly one facility (1d). However, by focusing solely on minimizing average distance, this approach can leave some community members unacceptably far from service, highlighting the need for models that explicitly consider equity.
2.1.1 Incorporating Equity into Facility Location Models
The goal is to replace the standard p-median objective function, given by (1a), with an alternative function. The new objective must adequately quantify the equitable access of our population without compromising computational scalability of the facility location model. Early attempts at this focused on protecting the worst-off residents. The traditional p-center model (Hakimi, 1965) minimizes the maximum distance that any resident must travel by introducing a single variable $z$ that captures the maximum distance:
@@ 2.1.1 Incorporating Equity into Facility Location Models
A natural extension is to minimize a convex combination of the two objective functions. Let $\gamma\in[0,1]$, then this appropriately termed “p-centdian” model (Halpern, 1976) may be written as:
(p-Ctdn)minimize $\gamma z+(1-\gamma)\frac{1}{T}\sum_{r\in R}\sum_{s\in S}p_{r}d_{r,s}y_{r,s},$
[…]
The value of $\gamma$ allows the user to weight the average distance term and the maximum distance term in the objective function. However, $\gamma$ is not readily interpretable beyond the obvious that larger values lead to the max term taking on greater importance. Moreover, due to the inclusion of this term, the p-centdian does not scale computationally as well as the p-median model. Along with the p-median and p-center models, we computationally explore the p-centdian model in detail in Section 5.
The Discrete Ordered Median Problem (DOMP) provides a general framework in which to view (p-Med), (p-Ctr), and (p-Ctdn). The objective function of the DOMP can be understood as minimizing a weighted sum of the distances from each resident to their closest open facility, i.e., for a resident located in area $r$, this distance is given by $\sum_{s\in S}d_{r,s}y_{r,s}$. The weight given to a particular resident is a penalty that depends on the position of $\sum_{s\in S}d_{r,s}y_{r,s}$ ordered relative to all other residents, and thus, accounts for equity. When all weights are equal, we obtain (p-Med), and at the other end of the spectrum, when all weights not corresponding to the resident traveling the furthest are set to zero, we obtain (p-Ctr).
Various linearizations have been developed which allow the DOMP to be formulated as a linear integer optimization problem (Nickel, 2001). Moreover, several custom algorithms have been proposed to solve the DOMP; leveraging techniques such as branch and cut (Boland et al., 2006; Marín et al., 2009; Marín, 2011; Deleplanque et al., 2020), branch and bound (Cherkesly et al., 2025), and Benders decomposition (Ljubić et al., 2024). For example, Ljubić et al. (2024) present a Benders decomposition method that matches or exceeds the current state-of-the-art on randomly generated instances of DOMP in computational tests. This method solved roughly 25% of instances of the largest problem size of $|R|=|S|=200$ within a two hour time limit. Through extensive testing, Cherkesly et al. (2025) demonstrate that their specialized branch-and-bound method successfully solves around 25% of instances of size $|R|=|S|=500$ within a two hour time limit. These include problems equivalent to (p-Med) and (p-Ctr) as well as more general DOMP problems. Here, our goal is to solve the equitable facility location problem on data sets with $|R|=2067$ and $|S|=350$ on average, and as large as $|R|=30,094$ and $|S|=8,274$. In order for the methodology to be accessible to a wide range of practitioners, we also wish to use an “out-of-the-box” integer programming solver, as opposed to a specialized method, therefore we seek an alternative framework to account for equity.
@@ 2.2 Measuring Inequality
Thus far, we have presented fundamental examples of facility location models focused on (i) efficiency, i.e., p-median; (ii) dispersion, i.e., p-center; and (iii) a parameter-controlled weighting of efficiency and dispersion (the p-centdian). In addition to these, many other metrics have been proposed to incorporate equity considerations in facility location, although each has important limitations. To enable us to consider the alternatives in context, we first discuss the properties required to appropriately measure equity.
@@ 2.2.1 Properties Required for Equity Assessment
The chosen location of services and amenities characterizes how a burden, namely travel distance, is distributed across a population. Thus, these decisions have significant equity implications and substantially impact community well-being, particularly for marginalized populations. For example, one key contributor to food insecurity is the travel distance required to reach a store offering affordable fresh produce, moreover, this burden is not equitably distributed in the majority of urban areas across the US (Neff et al., 2009).
The question of how best to quantify “equitable distribution” has long been considered in the environmental justice literature. More recently, clear criteria have been established for metrics used to evaluate distributions of both amenities and burdens; see Logan et al. (2021) and the references therein. There are six key properties that an inequality measure should possess if it is used to evaluate the distribution of a singular amenity or burden. For completeness, we note that a seventh and final property is included, however, this “multivariate” property is not relevant to our discussion and is omitted from the below.
@@
2.2.2 Review of Existing Measures (Suitability and Scalability)
@@
As we have seen, the maximum value metric, used in p-center models, optimizes the experience of the worst-off community member. In doing this, both the maximum value and range metrics suffer from considering only extreme values while ignoring the overall distribution. Moreover, neither satisfies the principle of transfers. Several other measures attempt to capture dispersion by examining deviation from the mean, including mean absolute deviation, variance, and the variance of logarithms. However, these also fail to satisfy the principle of transfers (Allison, 1978; Erkut, 1993).
@@
More sophisticated inequality measures have also been explored, including Schutz’s Index, the coefficient of variation, the Gini coefficient, Theil’s entropy coefficient, and Atkinson’s coefficient. While some of these metrics satisfy the transfers principle, they have other critical drawbacks. The Gini coefficient, despite its widespread use in facility location models (Mandell, 1991; Mulligan, 1991; Drezner et al., 2009; Lejeune and Prasad, 2013; Alem et al., 2022), measures only dispersion without considering the overall level of the distribution. Thus, it may prefer a solution where all residents are equally far from a facility over one with shorter but varied distances – clearly problematic for accessibility. The coefficient of variation and Theil’s coefficient suffer from the same limitation. This problem, i.e., violating scale dependence, is often overcome by augmenting the preferred measure of dispersion with an efficiency metric, a key example of this being the p-centdian objective function. However, it is often not clear how to “best” balance dispersion and efficiency metrics. In addition, the resulting measure is not guaranteed to satisfy all six of the properties necessary to assess equity.
Alternative approaches proposed to incorporate equity in facility location models include lexicographical optimization of multiple objectives (Ogryczak, 2009) and general p-norm optimization approaches (Karsu and Morton, 2015; Olivier et al., 2022; Gupta et al., 2023; Chen and Hooker, 2023). However, these methods either do not scale computationally, cannot evaluate both desirable and undesirable quantities, or lose their normative significance when applied to burdens rather than benefits (Barbati and Bruno, 2017).
@@
Equally-distributed equivalents (EDEs) offer a promising direction as they consider both the center and spread of a distribution in a single measure. The Atkinson EDE (Atkinson, 1970) has been applied to facility location, but makes the implicit assumption that higher values represent better welfare. Thus, it is inappropriate for measuring burdens such as travel distance because it does not satisfy the mirror property (Cox Jr., 2012). Furthermore, computational studies have shown that optimizing the Atkinson EDE becomes intractable for realistically-sized problems (Barbati and Bruno, 2017).
@@
Indeed, computational tractability emerges as a major barrier for many of the proposed equity metrics discussed. Recent comprehensive studies found that several measures are too complex to incorporate into optimization models that can handle real-world problem sizes (Barbati and Bruno, 2017; Chen and Hooker, 2023).
As a concrete example, we present Table 1 taken directly from Barbati and Bruno (2017), and adapted only for the purposes of consistent notation. Observe that in the largest instances with $|R|=|S|=60$ and $k=10$, only two of the ten measures considered could be solved to optimality within the 2 hour time limit for all instances. Recall that the problems we are interested in are orders of magnitude larger; hence the computational results suggest that we will be unable to optimize the majority of the measures, either directly or when combined with an efficiency measure.
Measure $|R|=|S|=20$ $|R|=|S|=40$ $|R|=|S|=60$
$k=2$ $k=10$ $k=2$ $k=10$ $k=2$ $k=10$
Center (Maximum) <1<1 <1<1 16 4 43 20
Range 1 4 17 21 464 269
Mean Absolute Deviation <1<1 <1<1 24 264 166 N/A
Variance 3 <1<1 205 810 1206 N/A
Maximum Absolute Deviation 3 <1<1 7 6 143 N/A
Absolute Difference 6 2 46 2656 1123 N/A
Sum Maximum Difference Absolute 13 3 56 767 1004 N/A
Schutz’s Index 3 3 81 N/A 1033 N/A
Coefficient of Variation 4 22 130 N/A 1588 N/A
Gini Coefficient 5 7 182 N/A 4022 N/A
Table 1: (Barbati and Bruno, 2017). Average solve time (in seconds) over 100 randomly generated test problems with $|R|$ demand points and $|S|$ potential facility locations, selecting $k$ locations. “N/A” denotes that an optimal solution could not be found for at least one of the instances within the time limit of 2 hours. Computations were performed using the solver Cplex v12.00 on a Pentium IV with 2.40 GigaHertz and 4.00 GigaBytes of RAM.
@@
The lack of a metric that demonstrably satisfies all six properties necessary to assess equity, in addition to computational challenges, had left the field of environmental justice without consensus on an appropriate and practical approach to measure equity in facility location problems. This fact motivated the development of the Kolm-Pollak EDE, which uniquely satisfies all key properties required for equity assessment in environmental justice applications. However, like many previous measures, its nonlinear form has limited its use to descriptive analysis rather than prescriptive optimization. This research bridges this gap. By developing a computationally tractable approach to optimizing the Kolm-Pollak EDE, we enable its application to large-scale facility location problems, allowing transportation planners to not just measure inequity but actively design for equity.